The area of a parallelogram is given by the formula Upper A equals bh. The area of the parallelogram on the right is 126 sq. units. Find its base and height.
b=4x+1
h=x+1

Answer by ikleyn(52803)

.
First step:  solve this quadratic equation  (4x+1)*(x+1) = 126  and find  x.


Second step:  having "x" found, calculate h = 4x+1  and  b = x+1.


At this point I completed my duty as a tutor.  The rest is on you.


Do not forget to post your  "Thanks"  to me for my instructions,  when you complete calculations.